# GOFA: A Generative One-For-All Model for Joint Graph Language Modeling

### v1

@@ 3.3 Unified Task Representation in GOFA
The formulation in Equation 1 provides a natural way for users to query the graph by selecting a NOG. However, tasks in the graph domain have different levels, such as node-, link-, and graph-levels, and different tasks have distinct objectives and target nodes. To cover these tasks under the same formulation, inspired by OFA Liu et al. (2023a), we convert all tasks into subgraph tasks and connect all target nodes in a graph to a virtual prompt node as the NOG for generation (i.e., $P$ nodes in Figure 1). Specifically, for node- and link-level tasks, we sample a $k$-hop rooted subgraph surrounding the target node or node pairs for node- or link-level tasks and connect the prompt NOG node to target nodes. For graph-level tasks, the prompt NOG node will connect to all nodes in the graph. This design has several advantages: (1) We can specify the prompt node as the NOG input to the LLM decoder for generative modeling. Therefore, the distribution of all tasks can be unified into a single space; (2) The text feature for the prompt node describes the task details. Connecting the prompt node to target nodes helps the prompt node query the most important knowledge from the input graph through attention and message passing. This encourages the model to dynamically learn from the message-passing process subject to prompt instead of only statically learning from the node text $x(v)$.
@@ 3.4 Large-Scale Pre-training
In this section, we describe the self-supervised pre-training of GOFA . The goal of the pre-training is to let the GOFA model obtain the ability to query graph structure and context but retain the ability to reason about plain text. Specifically, we perform the pre-training task using MAG240M Hu et al. (2021a) and Ultrachat200k Ding et al. (2023) datasets. Details about the datasets can be found in Appendix C. For each node $v$ in the MAG240M, we sample an ego-subgraph $G$ around $v$ to form a training example. We randomly generate a unique node ID (such as [Node A]) for each node in the training sample and append it to the original node text. This ID will serve as a basis for querying nodes in the graph. We design three pre-training tasks: sentence completion, structural tasks on MAG240M, and question-answering tasks on Ultrachat200k. Figure 1 shows an example of each task. We describe the details of each task below.
Sentence completion task. The objective of the sentence completion task is to train the model to reason about the rest of the text in a node given both the existing text and the information in the rest of the graph. First, we assign the root node as the target for each sampled training graph. Next, we randomly select $n$ nodes in the graph and the target node. All selected nodes’ texts are split into halves. The first half forms node text $x(v)$, and the second half becomes the target $y$ to generate. Additionally, we connect a prompt node $v_{p}$ for each selected node, as shown in Figure 1. Note that for all prompt nodes, we use directed edges to connect them with nodes in the input graph, such that only the information in the input graph will be passed to the prompt node for decoding, and information in different prompt nodes does not influence each other. The text in each prompt node asks the model to complete the sentence of the selected node it connects. This sentence-completion pre-training task adapts LLMs’ standard ability to the graph context.
Structural tasks. The objective of the structural tasks is to pre-train GOFA to understand basic graph structural properties. In light of this, we design the shortest paths and common neighbors reasoning tasks between nodes. Specifically, For each training subgraph sample, we randomly sample $n$ nodes as the selected nodes. For each selected node, we ask the model to compute the shortest path distance between it and the root node and output all possible shortest paths between them. Meanwhile, we also ask the model to output the number of common neighbors the selected node and the target node have and the node IDs of their common neighbors. The prompt node $v_{p}$ will connect to both the target and the selected nodes since our structural tasks need the model to reason about two nodes simultaneously. The text in the prompt node will be the corresponding question. Through these two tasks, the model is expected to gain the ability to identify basic graph structures, which are critical to downstream tasks.
Question answering task. Unlike language corpus, which naturally contains many question-and-answer (QA) pairs, graph data usually only contain objective descriptions of entities. Nevertheless, for the model to be fluid in tasks, we need the model to understand user prompts and be sensitive to different tasks. Hence, we synthesize a QA-chain dataset from Ultrachat200k, as shown in Figure 1. A language QA sequence is converted to a chain where nodes with question texts alternate with nodes with answer texts, which are connected by directed edges to represent the conversation order. The last question becomes the text on prompt node $v_{p}$, which is connected to every node in the chain, and the last answer is the target text $y$ (see Figure 1 QA-Chain Task for an example). This QA task provides QA pseudo-graphs missing from the common graph corpus, and we found it critical for allowing the model to be responsive to arbitrary tasks expressed in free-form questions.
@@ 4 Related work
Prompt Learning in Graph: The success of foundation models inspired many works that aim to adapt their power to the graph domain. Earlier attempts designed a graph prompting mechanism such that a trained model can adapt to new data by fine-tuning a soft prompting vector Liu et al. (2023b); Yu et al. (2023); Sun et al. (2023); Xia et al. (2024). GraphPrompt Liu et al. (2023b); Yu et al. (2023) pretrains on link prediction tasks, and then finetune a prompt matrix for each downstream node or graph classification task. All in One Sun et al. (2023) designs prompt tokens that are used to modify node features and then take a meta-learning paradigm for multi-task learning. Subsequent works extend graph prompts to allow in-context learning without weight update Huang et al. (2023a); Galkin et al. (2023). For example, Prodigy Huang et al. (2023a) designs a few-shot prompting graph structure connecting label nodes and data nodes related to support set samples, and trains a GNN for link prediction between label and data nodes. However, these works only tackle limited types of tasks and do not generalize to new domains. Hence, researchers propose integrating LLM into the graph learning and benefit from LLMs’ versatility.
@@
LLMs as enhancers: One direction uses LLMs to convert the text features of graphs to unified representations Liu et al. (2023a); Chen et al. (2023); Li et al. (2024); He et al. (2024a) for downstream graph models to distinguish and transfer knowledge between different domains. For example, OFA Liu et al. (2023a) uses LLM to unify the input features in different datasets and transforms multiple types of graph classification tasks into a unified binary classification format, which allows a single graph model to be trained on diverse tasks and domains. However, OFA could not handle regression tasks. TAPE He et al. (2024a) utilizes LLM to generate question answers and explanations as enhanced node features and use them as input to train a GNN for node classification tasks. Such approaches have good structural modeling ability because of the explicit use of graph structure. However, they usually cannot generate free-form output to handle arbitrary tasks.
@@
LLMs as predictors: Another line of research proposes using LLMs as predictors and aligning graph representation with LLM inputs. Preliminary attempts flatten graphs into text representations and feed them into LLM Chen et al. (2023); Zhao et al. (2023b); Guo et al. (2023). These approaches can benefit from LLM for task fluidity but fail to model structural information unique to graph data properly Zhao et al. (2023b); Mao et al. (2024); Ye et al. (2023). Alternatively, in molecule modeling, LLM is directly fine-tuned on molecules’ SMILE string representations Zhao et al. (2023a); Qian et al. (2023). However, the application for such models is narrow to molecule and chemistry. Realizing this problem, follow-up work extends methods in vision-language domain Alayrac et al. (2022); Li et al. (2023) to the graph domain and train adapters to link graph model outputs to LLM Tang et al. (2023, 2024); Huang et al. (2024); Zhang et al. (2024); He and Hooi (2024). For example, GraphGPT Tang et al. (2023) first implements a text-structure alignment between graph representation and text embedding to pretrain a GNN, then takes GNN output to insert to language space and performs instruction tuning for downstream tasks. Inspired by Q-former Li et al. (2023), GraphTranslator Zhang et al. (2024) aligns node and text tokens from pre-trained GNN and LLM. UniGraph He and Hooi (2024) pretrains GNN using masked word prediction and then tuning a projector to map graph embedding to language space and enable zero-shot learning. However, the GNN and LLM parts of these methods are usually detached, meaning the prompt information can not attend to the message-passing process. More related work about GNNs and transformers is included in Appendix D.
@@ 5 Experiment
This section evaluates the proposed methods by answering the following four questions: Q1: Can GOFA use graph information to better model TAG and help complete the target sentence? Q2: Does the pre-trained GOFA help with zero-shot learning? Q3: Which portion of GOFA gives it unique TAG modeling and in-context learning ability? Q4: Can we transfer the prior knowledge in the pre-trained GOFA to downstream fine-tuning tasks?
@@ 5.1 GOFA pre-training
To answer Q1, we pre-train two GOFA models using ICAE models on Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023), optimizing the objective in Equation 4. The pre-training datasets include MAG240M Hu et al. (2021a), UltraChat200k Ding et al. (2023), and our proposed structural learning tasks as introduced in Section 3.4. The training details can be found in Appendix F. Figure 2 shows a training loss decrease for the number
@@
Llama2-7B 14.31 1.627 2.162
Mistral-7B 13.03 0.859 1.455
GOFA-Llama 4.844 0.580 0.746
GOFA-Mistral 4.811 0.553 0.531
Table 2: GOFA training loss.
of seen tokens. We report the perplexity in Table 2 and compare GOFA with base LLM finetuned on the MAG240M dataset. Note that during pre-training, we only update the weight of the GNN layers, and GOFA ’s lower perplexity shows that the structural and semantic information in the node’s neighbor can effectively help complete the sentence with more relevance than the original LLM. We observe that fine-tuned LLM does not improve much from the base LLM as it is already well-trained on large-scale academic datasets and cannot directly use the neighbor information to refine the generation. Besides sentence completion, another important GOFA pre-training objective is the structure learning ability; we report shortest path distance and common neighbor count prediction results in Table 2, compared with LLM models whose inputs are textualized graphs, with descriptions of edge connections. We see a significant performance improvement of GOFA over base LLM, showing that a difficult graph task for LLM can be well solved by the GNN layers with better structure modeling ability.
@@ 5.2 Zero-shot learning with GOFA
Table 3: PubMed link instruction tuning on zero-shot setting (Acc).
Task Cora-Link
Way 2
Llama2-7B 49.72
Mistral-7B 41.34
OFA-Llama2 52.22
GOFA-Llama2-Pubmed 62.36
GOFA-Mistral-Pubmed 65.91
To answer Q2, we performed zero-shot experiments on various graph tasks. Despite using QA-chain data in the pre-training stage, the graph data does not include knowledge about task formats like classification and does not output exact matches to the answers. Hence, we first perform instruction tuning on the pre-trained GOFA in Section 5.1 using two question-template datasets, arxiv, and Pubmed-link. We only use a part of the data, and the goal of instruction fine-tuning is not to let the model learn particular knowledge from these datasets but to make the model understand the task format described in Appendix F.3. We intentionally chose these datasets because the distribution is similar to the pre-training MAG240M dataset, which is important for fair zero-shot experiments.
@@
Table 4: Zero-shot experiment results with Arxiv instruction tuning (Accuracy). Bold and underlined shows best and runner-up results. L2 and M represent Llama-2 and Mistral LLM, respectively.
Task Cora-Node PubMed-Node WikiCS Products ExplaGraphs FB15K237 SceneGraphs
Way 7 3 10 10 2 10 QA
Llama2-7B 29.69 60.95 32.56 50.69 59.02 27.66 38.62
Mistral-7B 54.79 71.02 58.83 61.99 73.03 63.85 45.95
OFA-Llama2 27.70 56.42 18.5 - - - -
GraphGPT 18.13 70.11 - - - - -
UniGraph 69.53 72.48 43.45 66.07 - - -
GOFA-L2-arxiv-10K 63.56 65.26 47.98 36.85 53.61 28.47 23.52
GOFA-M-arxiv-10K 65.15 64.37 68.19 72.60 78.21 45.81 32.44
GOFA-M-arxiv-40K 71.20 73.11 70.49 75.83 79.56 55.96 33.06
@@
We perform 10-way classification fine-tuning on only 10k/40k arxiv data for one epoch and show zero-shot results in Table 4. While arxiv only contains node classification tasks, we observe that GOFA achieves very non-trivial performance on all node-level (Cora, Pubmed, WikiCS, Products), link-level (FB15K237), and graph-level (ExplaGraphs, SceneGraphs) tasks. GOFA also generalizes to different ways and even question-answering (SceneGraphs) tasks, showing its desirable fluidity. GOFA outperforms LLM and graph foundation model baselines on most datasets, and exceeds second best results by a large margin ($>5\%$) on WikiCS, Products, and ExplaGraphs, showing GOFA ’s ability to combine the advantage of both LLM and graph models. We also notice that GOFA does not perform as outstanding on the FB15K237 knowledge graph dataset, potentially because it is a link prediction task, which requires task-specific knowledge different from our arxiv instruction tuning data. Using a knowledge graph for tuning can potentially improve the performance and we leave this to future work. We also observe that GOFA is only achieving comparable performance to LLM on the SceneGraph dataset. We notice that SceneGraph contains object coordinate description which makes it easy for LLM to answer tasks about relative location. Whereas, our compressed representation might still discard relevant number information compared to directly prompting this information to LLM. Hence, the advantages of GOFA are not as apparent on tasks where text-only information is sufficient. We discuss potential solutions in Appendix F and leave this to future work.
5.3 Investigation of GOFA architecture
Figure 4: Ablation study on prompt edge.
Single Edge Double Edge
Cora-Link 49.6 59.4
Cora-Node 58.6 60.2
Products 2.8 22.8
WN18RR 6.4 14.6
Figure 5: Representation changed ratio.
For Q3, we observe one unique feature of GOFA is that its GNN layers are embedded into the LLM. A natural question is how important the interleaved GNN layers affect the LLM layer outputs. Hence, we compute a layer-wise Representation Change Ratio value, $\delta=||H^{t}-Q^{t}||/||Q^{t}||$ ($H^{t}$ undergoes MLP, residual, and gating modules), quantifying the change that GNN brings to the LLM outputs; a larger value means a higher dependence on graph information. We take the mean ratio of 100 data randomly sampled from each zero-shot dataset. The results in Figure 5 show that almost all GNN layers have a significant impact on the LLM output in both pre-trained and node(arxiv)/link(Pubmed-link) instruction-tuned versions. The pre-trained and arxiv versions have similar importance for each layer as they are all academic datasets and node-level tasks usually obtain enough information for neighbors within 2-hops. The link-finetuned version has larger third and fourth-layer importance, showing that the link task usually requires the model to consider a larger neighborhood to understand the graph structure.
Another important design in GOFA that enables in-context zero-shot learning is the injected prompt node with the user instruction. This allows the GNN to pass messages differently when users provide different prompts. We present a zero-shot ablation study on the connection between target and prompt nodes in Table 5. Single edge means that only the prompt node receives a message from the target nodes, and double edge means that target nodes also receive a message from the prompt node. The former results in static message passing independent of the prompt, and the latter in dynamic message passing. We simultaneously train GOFA -llama on Pubmed-link and arxiv datasets and evaluate their zero-shot performance. Without dynamic message passing, the fine-tuned model can only perform well on one dataset and fail to use its inductive knowledge when the task (Cora-link v.s. Cora-node) or the domain (Cora v.s. Products) shifts. This shows the importance of the prompt node design, which enables dynamic message-passing that can better use the pre-trained knowledge of the GNN and LLM compressor for zero-shot tasks. We answer Q4 with supervised training in Appendix B.

### v2

@@ 3.3 Unified Task Representation in GOFA
The formulation in Equation (1) provides a natural way for users to query the graph by selecting a NOG. Following OFA (Liu et al., 2023a), we convert all tasks into tasks on $k$-hop rooted subgraphs extracted around the target nodes. For node-level tasks, the target node is a single node in the graph. For link-level tasks, the target nodes are the node pair. If the target node is not specified (e.g., the task is a graph task), we set the default target nodes to all nodes in the graph. We connect a prompt node with the user query as NOG to all target nodes. GOFA completes the prompted input TAG by answering the query on the NOG, which still aligns with the proposed generative modeling framework. This design has several advantages: (1) All tasks are represented by a NOG, so the distribution of all tasks can be unified into a single space, helping the model generalize to unseen tasks from learned task representations; (2) The text feature for the prompt node describes the task details. Connecting the prompt node to target nodes enables the prompt node to query the most important knowledge from the input graph through attention. This ensures the output embedding for NOG is conditionally learned from the GNN process subject to the different prompts. Conversely, most of the previous works (He & Hooi, 2024; Tang et al., 2023; 2024; Zhang et al., 2024) only computed a fixed embedding for each node before any prompt is introduced.
@@ 3.4 Large-Scale Pre-training
As discussed in Section 2 and Section 3.1, we design self-supervised pre-training tasks based on the three GFM properties to train GOFA. The training datasets include MAG240M (Hu et al., 2021a), Pubmed and Arxiv (Hu et al., 2021b) for academic knowledge, Wikikg90mv2 (Hu et al., 2021a) and WikiGraph (proposed by us) for semantic diversity, and Ultrachat200k (Ding et al., 2023) dataset for question-answering ability. Details about the datasets can be found in Appendix C. Each node is assigned a unique ID (e.g., [Node A]) to enable node querying in the graph. We design four pre-training tasks as shown in Figure 1. We describe the rationale of each task below and leave some implementation details and additional discussion in Appendix F and Appendix E.3.
Sentence Completion Task. This task aims for large-scale pre-training (GFM property one) by training GOFA to predict the remaining text in a node based on both the existing node text and the surrounding graph information. Such a task can be applied to any TAG without labeling, thus facilitating large-scale pre-training for GOFA to acquire diverse knowledge.
Structural Understanding Task. This task aims to provide structural modeling ability for GOFA (GFM property three). The structural task connects NOG randomly selected node pairs to generate the actual shortest path or common neighbors between them. Through these two tasks, the model is expected to gain the ability to identify basic graph structures fundamental for graph-related problems.
Question Answering Task. This task aims to ensure fluidity in generation for GOFA (GFM property two). Unlike language corpus, which naturally contains many question-and-answer (QA) pairs, graph data usually only contain objective descriptions of entities. Hence, we convert natural language Question-Answer sequences into chain graphs and connect a NOG with a question to the chain graph for open-ended answer generation. This essential task enables GOFA to be responsive to arbitrary downstream applications expressed in free-form text questions.
Information Retrieval Task. In most downstream tasks, GOFA links a prompt node to target nodes in the graph to address related problems. To facilitate effective information extraction, we design an information retrieval task where a NOG queries a target node using its node ID. The model must retrieve and isolate information specific to the queried node from the remaining target nodes, encouraging a message-passing process conditioned on the input, as discussed in Section 3.2.
@@ 4 Related work
Here we mainly discuss the two tracks of general graph models, and leave discussion about graph prompt learning and graph neural networks to Appendix D.
@@
LLMs as enhancers: One direction uses LLMs to convert the text features of graphs to unified representations (Liu et al., 2023a; Chen et al., 2023; Li et al., 2024; He et al., 2024a; Plenz & Frank, 2024) for downstream graph models to distinguish and transfer knowledge between different domains. For example, OFA (Liu et al., 2023a) uses LLM to unify the input features in different datasets and transforms multiple types of graph classification tasks into a unified binary classification format. TAPE (He et al., 2024a) utilizes LLM to generate question answers and explanations as enhanced node features. Such approaches have good structural modeling ability, but they usually cannot generate free-form output to handle arbitrary tasks.
@@
LLMs as predictors: Another line of research proposes using LLMs as predictors and aligning graph representation with LLM inputs. Preliminary attempts flatten graphs into text representations and feed them into LLM (Chen et al., 2023; Zhao et al., 2023b; Guo et al., 2023; Zhao et al., 2023a; Qian et al., 2023). These approaches can benefit from LLM for task fluidity but fail to model structural information unique to graph data properly (Zhao et al., 2023b; Mao et al., 2024; Ye et al., 2023). Realizing this problem, follow-up work extends methods in vision-language domain (Alayrac et al., 2022; Li et al., 2023) to the graph domain and train adapters to link graph model outputs to LLM (Tang et al., 2023; 2024; Huang et al., 2024; Zhang et al., 2024; He & Hooi, 2024). For example, GraphGPT (Tang et al., 2023) first implements a text-structure alignment between graph representation and text embedding to pretrain a GNN. LLaGA (Chen et al., 2024) creatively uses a template to represent a subgraph with pooled node embeddings for LLM input. Inspired by Q-former (Li et al., 2023), GraphTranslator (Zhang et al., 2024) aligns node and text tokens from pre-trained GNN and LLM. UniGraph (He & Hooi, 2024) pretrains GNN using masked word prediction and then tuning a projector to map graph embedding to language space and enable zero-shot learning. However, the GNN and LLM parts of these methods are usually detached, meaning the prompt information can not attend to the message-passing process.
@@ 5 Experiment
This section evaluates the proposed methods by answering the following four questions: Q1: Are the pre-training tasks in GOFA effective for graph-language modeling and structure understanding? Q2: Does the pre-trained GOFA help with critical general graph model application, zero-shot learning? Q3: Is using GOFA more advantageous than LLMs in graph tasks? Q4: Does GOFA have the fluidity to handle open-ended graph-related tasks? Additionally, we also include supervised experiments in Appendix F.5.
@@ 5.1 GOFA pre-training
Mistral-7B 30.12 1.254 1.035
GOFA-SN 26.20 - -
GOFA 21.34 0.634 0.326
To answer Q1, we pre-train the GOFA model using ICAE models on Mistral-7B (Jiang et al., 2023), optimizing the objective in Equation (1) using the proposed tasks. The training details can be found in Appendix F.3. After training, we evaluate the perplexity of both GOFA and base LLM on Cora, Product, and Wikics datasets (all three are not included in the pre-training). We report the perplexity in Table 1. Note that during pre-training, we only update the weight of the GNN layers, and GOFA ’s lower perplexity shows that the structural and semantic information in the node’s neighbor can effectively help complete the sentence with more relevance than the original LLM. Further, to validate that training of GOFA will not affect the original LLMs’ ability, we input GOFA with single node graphs without any connections (denoted as GOFA-SN) to evaluate the perplexity, as shown in Table 1. We can see that without connection information around the center node, generation on a single node graph remains comparable to LLM and even better due to the pre-training process, showing that GOFA training does not destroy the desirable property of a pre-trained LLM. Besides sentence completion, another important GOFA pre-training objective is the structure learning ability. We report the shortest path distance and common neighbor count prediction results in Table 1, compared with LLM models whose inputs are textualized graphs with descriptions of edge connections. The datasets we used were Cora and Product. We see a significant performance improvement of GOFA over base LLM, showing that a difficult graph task for LLM can be well solved by the GNN layers with better structure modeling ability.
@@ 5.2 Zero-shot learning with GOFA
Table 2: Zero-shot experiment results with instruction tuning (Accuracy).
Task Cora-Node WikiCS Products ExplaGraphs Cora-Link
Way 7 2 10 5 47 10 5 2 2
LLama2-7B 47.92 73.45 40.10 58.77 27.65 58.71 64.33 57.76 48.15
Mistral-7B 60.54 88.39 63.63 71.90 43.99 70.16 74.94 68.77 49.43
OFA-Llama2 28.65 56.92 21.20 35.15 19.37 30.43 39.31 51.36 52.22
GraphGPT 44.65 - - - 18.84 - - - 50.74
UniGraph 69.53 89.74 43.45 60.23 38.45 66.07 75.73 - -
ZeroG 64.21 87.83 31.26 48.25 31.24 51.24 71.29 - -
LLaGA 51.85 62.73 - - 23.10 34.15 39.72 - 88.09
GOFA-T 70.81 85.73 71.17 80.93 54.60 79.33 87.13 79.49 85.10
GOFA-F 69.41 87.52 68.84 80.62 56.13 80.03 88.34 71.34 86.31
Table 3: Zero-shot experiment results with instruction tuning on FB15K237 and Scene Graphs.
Task FB15K237 SceneGraphs
Format 10-Way QA
Llama2-7B 48.32 38.62
Mistral-7B 62.48 45.95
GOFA-T 73.59 34.06
GOFA-F 80.69 31.36
Figure 4: Performance vs pre-training sample size.
Figure 5: Pre-training Tasks Ablation Study.
To answer Q2, we performed zero-shot experiments on various graph tasks. Despite using QA-chain data in the pre-training stage, the graph data does not include knowledge about task formats like classification and does not output exact matches to the answers. Hence, we first instruction-tuned the pre-trained GOFA in Section 5.1 on a small amount of data. We report the zero-shot results of two GOFA instruction tuning settings named GOFA-T and GOFA-F, as shown in Table 2 and Table 3. GOFA-T includes node and link classification tasks from Arxiv and Pubmed and GOFA-F addtionally adds MAG240M and Wiki90mv2 datasets. The instruction-tuning details can be found in Appendix F.4. Note that the zero-shot datasets are unseen during both pre-training and instruction finetuning. The goal of instruction fine-tuning is not to let the model learn particular knowledge from these datasets but to make the model understand the task format described in Appendix F.4.
@@
While the instruction-tuning dataset only covers the relatively small spectrum of the graph datasets, we observe that GOFA achieves very non-trivial performance on all node-level (Cora-Node, WikiCS, Products), link-level (FB15K237, Cora-Link), and graph-level (ExplaGraphs, SceneGraphs) tasks. GOFA also generalizes to different ways and even question-answering (SceneGraphs) tasks, showing its desirable fluidity. GOFA outperforms LLM and graph foundation model baselines on most datasets and exceeds best baselines by a large margin ($>10\%$) on WikiCS, Products, FB15K237 and ExplaGraphs, showing GOFA’s ability to combine the advantage of both LLM and graph models. GOFA not only achieves remarkable results on the knowledge graph and academic graph, which are proximal to the trained data but also excels in Products and ExplaGraphs whose distribution shifts significantly from training data, which further highlights GOFA ’s substantial generalizability. Meanwhile, we observe that GOFA is only achieving comparable performance to LLM on the SceneGraph dataset. We suspect that the instruct-tuning data contains information-dense texts, reducing the model’s ability on common sense questions that this dataset requires. In the future, we plan to diversify instruction-tuning datasets with common sense knowledge to enhance such ability.
We further conducted the same experiments on intermediate pre-training checkpoints, and show results in Figure 4. We observe that as the model witnesses more pre-training samples/tokens, the downstream task performance also increases significantly, confirming the importance of large-scale pre-training on graph data. The performance continues to improve, meaning that the model can potentially scale to higher capability with more samples; we leave this to future work. In Figure 5, we plot the instruction-tuning performance when we remove the Wikipedia datasets and information retrieval task (w/o R+W), only remove the retrieval task, (w/o R), and full tasks. We can see that Wikipedia datasets improve the model performance of all the datasets for the diverse corpus it introduced. The retrieval tasks particularly improve the knowledge graph performance due to the improved ability to retrieve key correlations between target entities. These show the necessity and effectiveness of the overall pre-training task selection and design. In Appendix B.1, we further conduct zero-shot experiments on molecule datasets.
5.3 Comparing GOFA with LLMs
Table 4: Comparison between GOFA and LLM with the same input.
Task ExplaGraphs Time WikiCS Time Cora-Link Time FB15k237 Time
Metric Acc $\uparrow$ sec/sample $\downarrow$ Acc $\uparrow$ sec/sample $\downarrow$ Acc $\uparrow$ sec/sample $\downarrow$ Acc $\uparrow$ sec/sample $\downarrow$
LLM-N 74.13 1.50 OOM OOM 50.36 3.84 51.25 3.92
GOFA-F 79.49 0.48 71.17 2.43 85.10 1.67 73.49 3.37
Improvement 7.23% 68.00% NA NA 68.98% 56.51% 43.40% 14.03%
Answering Q3 is critical to understanding the necessity of the GNN layers and the effectiveness of GOFA as a general graph model. We compare GOFA to LLM whose textual prompt contains the same information as the input graph to GOFA. Specifically, for a GOFA input graph, we concatenate all node texts as the prompt and append the connection information to it, as in "Node A connects to Node B". The text is then combined with task and question descriptions as input to an LLM for classification tasks. Approaches similar to this are widely adopted and acknowledged (Chen et al., 2023; Fatemi et al., ). We present both the classification performance and per sample inference time in Table 4 and denote the LLM method as LLM-N. We observe impressive performance improvement of GOFA on all datasets, even when the LLM is prompted with the same information, showing that GOFA, with the help of the GNN and interleaving design, utilizes the graph information much more effectively. Moreover, we also observe a fundamental reduction in inference costs, confirming our analysis in Section 3.2 that, with the same input, GOFA is more efficient than LLMs. Note that when the input size is large, such as in WikiCS, LLM struggles with high memory consumption of the long sequence, whereas the GOFA avoids that by leveraging the sparsity of graph data and using edge information to compute the most important attention information.
5.4 GOFA Responses on Diverse Tasks
Finally, we answer Q4 by providing generation examples of GOFA in Figure 6, where we prompt the same citation graph differently and achieved corresponding and high-quality responses. The top and middle examples have the same connection for their NOGs (both connected to the same five nodes), but when we change the prompt text on the NOGs, the generated texts also adjust accordingly, utilizing the neighbor node information, validating that the message-passing is conditioned on the prompt. As in the bottom example, we can also prompt the graph differently by connecting the NOG to two target nodes and querying about the shortest path distance. In this case, the model successfully generates actual paths between the two nodes, which is an ability not seen in traditional graph models that can only output numerical predictions about the path length. These examples demonstrate GOFA’s outstanding ability to answer open-ended questions. More examples are provided in B.4.
Figure 6: GOFA diverse responses to open-ended questions.